# A Survey of Multimodal Large Language Model from A Data-centric Perspective

### v1

@@ Abstract.
Human beings perceive the world through diverse senses such as sight, smell, hearing, and touch. Similarly, multimodal large language models (MLLMs) enhance the capabilities of traditional large language models by integrating and processing data from multiple modalities including text, vision, audio, video, and 3D environments. Data plays a pivotal role in the development and refinement of these models. In this survey, we comprehensively review the literature on MLLMs from a data-centric perspective. Specifically, we explore methods for preparing multimodal data during the pretraining and adaptation phases of MLLMs. Additionally, we analyze the evaluation methods for datasets and review benchmarks for evaluating MLLMs. Our survey also outlines potential future research directions. This work aims to provide researchers with a detailed understanding of the data-driven aspects of MLLMs, fostering further exploration and innovation in this field.
Multimodal Large Language Models; Data-centric Artificial Intelligence
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACM
@@
In recent years, we have witnessed rapid advancement of large language models (LLMs) and multimodal large language models (MLLMs) (Zhao et al., 2023d; Wu et al., 2023b). MLLMs, such as GPT-4 (OpenAI, 2023), Flamingo (Alayrac et al., 2022), BLIP2 (Li et al., 2023f) and X-InstructBLIP (Panagopoulou et al., 2023), integrate multiple modality information, demonstrating impressive comprehension and generation capabilities. These models achieve competitive performance in traditional multimodal tasks, such as visual recognition (Zhang et al., 2024a), video understanding (Xu et al., 2021a; Tang et al., 2023), speech recognition (Min and Wang, 2023) and 3D understanding (Guo et al., 2023; Hong et al., 2023). Moreover, their excellent language understanding capacity enables strong performance in text-rich tasks, such as question answering (Hu et al., 2024), multi-dialog conversation and logical reasoning (Xu et al., 2024; Li et al., 2023l).
Most existing MLLMs focus on modifying model architecture to explore the use of information from multiple modality (Longpre et al., 2023; Xie et al., 2024a; Soboleva et al., 2023; Xie et al., 2024b; Kaddour, 2023; Ye et al., 2024). While model effectiveness is crucial, data also significantly impacts the success of MLLMs. For example, Hoffmann et al. (2022) demonstrate that in order to scale up models, it is necessary to increase the training data size. Beyond data quantity, data quality is equally important. Previous research (Sorscher et al., 2022) indicates that carefully curated datasets can enable smaller models to achieve comparable performance to larger ones. However, comprehensive studies on data curation and utilization for MLLMs are still lacking. Therefore, this study aims to provide a comprehensive understanding of MLLMs from a data-centric perspective.
In contrast to model-centric approaches that prioritize architectural enhancements while relying on fixed datasets, data-centric perspectives emphasize iterative improvements to datasets to enhance performance. Within the scope of data-centric MLLMs, our focus lies on leveraging the heterogeneous nature of data modalities, enhancing data structure, increasing data quantity, and elevating data quality to augment MLLMs (Zha et al., 2023). Our discussion answers three key questions from a data-centric perspective at different stages of MLLMs:
  • Q1: How to collect, select, and manage data for MLLMs? The substantial data quantity requirements and the heterogeneity of multimodal data pose challenges in gathering, selecting, and effectively managing data for model training. Different training stages of MLLMs also lead to varying data type requirements.
@@
Difference between this survey and the existing one. Several existing surveys have focused on LLMs (Zhao et al., 2023d; Hadi et al., 2023; Naveed et al., 2023) and MLLMs (Wu et al., 2023b; Zhang et al., 2024c) from a model-centric perspective, but they lack an in-depth analysis of data-centric aspects. Recently, some surveys have started to focus on data preparation for LLMs, such as data management methods (Wang et al., 2023b), data selection methods (Albalak et al., 2024), and comprehensive reviews of LLM datasets (Liu et al., 2024a). However, these surveys primarily focus on data management and selection methods for text-only LLMs and do not provide a thorough analysis of the data processing pipeline for MLLMs. Although Zhang et al. (2024c) summarize datasets for MLLMs, they fall short of providing a comprehensive analysis of these datasets. The most closely related work to ours is data-centric artificial intelligence (DCAI) (Zha et al., 2023; Jarrahi et al., 2022; Jakubik et al., 2022; Polyzotis and Zaharia, 2021; Whang et al., 2023), which also focuses on data-centric views of AI research but does not specifically analyze LLMs and MLLMs.
@@
Contributions. In this survey, we reviewed the literature on the advances of MLLMs from data-centric perspectives. We provide researchers and developers with a general and comprehensive understanding of the latest developments in the field of MLLM data side. The key contributions of this survey are summarized as follows:
@@
  • Data evaluation benchmarks. We outline commonly used evaluation benchmarks from a data-centric perspective.
@@
The remainder of this article is organized as follows: Section 2 introduces the preliminaries for LLMs and MLLMs, and discusses the motivations for analyzing them from a data-centric perspective. Sections 3 to 5 summarize the main stages of collecting, processing and selecting data for MLLMs training. Section 6 summarizes data evaluation methods and existing evaluation datasets for MLLMs according to evaluation tasks. Section 7 discusses open issues and highlights several future research directions in this field. Finally, we conclude the survey in Section 8. Our Github repository can be found in https://github.com/beccabai/Data-centric_multimodal_LLM.
@@ 2.1. Large Language Models
A significant breakthrough came with the introduction of pre-trained language models, which aimed to capture general language representations by training on vast amounts of unlabeled text data. Models such as BERT (Kenton and Toutanova, 2019) and GPT (Radford et al., [n. d.]) demonstrated the effectiveness of pre-training and fine-tuning for various NLP tasks. Building upon these advancements, LLMs have taken the concept of pre-training to an unprecedented scale, with models containing billions of parameters and trained on massive corpora spanning hundreds of gigabytes to terabytes of text data (Brown et al., 2020; Raffel et al., 2020)
LLMs exhibit several notable properties that distinguish them from previous language models. One fundamental characteristic is the scaling law, which describes how the performance of large language models improves as they scale in terms of model size, training data, and computational resources. These laws provide insights into the relationships between the number of parameters, the amount of training data, the computational power used, and the resulting model performance. Empirical evidence has shown that the performance of LLMs on downstream tasks improves as the model size and the amount of pre-training data increase (Kaplan et al., 2020; Hoffmann et al., 2022). This power-law relationship suggests that larger models trained on more data tend to capture more complex patterns and generalize better to new tasks.
@@
Another intriguing property of LLMs is the emergence of abilities that were not explicitly trained for, often referred to as emergent abilities (Wei et al., 2022). This ability suggests that LLMs can capture and leverage complex linguistic patterns and knowledge from the pre-training data, enabling them to perform tasks beyond their original training objective. Furthermore, LLMs have demonstrated the capability of in-context learning (Brown et al., 2020), where they can perform tasks based on a few examples provided in the input prompt, without the need for explicit fine-tuning. This highlights the models’ ability to rapidly adapt to new tasks and generalize from limited examples.
@@
The development and deployment of LLMs typically involve three key stages: training, adaptation, and evaluation (Zhao et al., 2023d). The training stage focuses on learning general language representations from large-scale unlabeled corpora, capturing the underlying patterns and structures of natural language. The adaptation stage involves tailoring the pre-trained model to specific domains or tasks, often through techniques such as domain adaptation or task-specific fine-tuning (Raffel et al., 2020). This step is crucial for optimizing the model’s performance on the target application and ensuring its outputs are aligned with the desired objectives. Finally, the evaluation stage involves assessing the performance of the adapted model through various metrics and benchmarks to ensure that it meets the required standards and performs effectively in real-world scenarios.
From a data-centric perspective, the performance and generalization capabilities of large language models (LLMs) are highly dependent on the quality and diversity of the data used throughout the different stages of development (meta llama, 2024). During the training stage, the pre-training data is crucial in shaping the model’s language understanding and generation abilities. Efforts have been made to curate high-quality datasets (Gao et al., 2020) that cover a wide range of domains and languages. In the adaptation stage, domain-specific data is often used to fine-tune the model and enhance its performance on targeted tasks. The evaluation stage involves using real-world data for inference and assessment, underscoring the importance of data representativeness and robustness.
However, challenges persist in addressing data bias, privacy concerns, and the need for more diverse and representative datasets that capture the richness and nuances of human language across different domains and cultures. Overcoming these challenges requires ongoing research in data collection, curation, and augmentation techniques (Raffel et al., 2020; Perez et al., 2022).
The development of LLMs has revolutionized the field of NLP, enabling significant advances in language understanding and generation. The scaling law, emergent capabilities, and in-context learning capabilities of LLMs have opened up new avenues for research and application.
@@ 2.2. Multimodal Large Language Models
Multimodal large language models (MLLMs) extend the capabilities of traditional LLMs by leveraging their comprehension, generation, and task-solving abilities across different modalities, including natural language (text and audio) and visual information (videos, images, and 3D models). By integrating LLMs with multimodal projectors, MLLMs can process various types of information, demonstrating strong understanding and analytical capabilities to address downstream tasks across various modalities (Wu et al., 2023b).
@@
The development of MLLMs is driven by the need to address real-world problems that often involve multiple modalities. For instance, in robotics, an MLLM can process visual input from cameras, interpret natural language instructions, and analyze structured data from sensors to perform complex tasks (Zeng et al., 2023). Similarly, in healthcare, MLLMs can analyze medical images, process electronic health records, and interpret patient-doctor conversations to aid in diagnosis and treatment planning.
The training of MLLMs typically involves three distinct stages: pre-training, adaptation, and evaluation. First, the pre-training stage involves training the model on large-scale multimodal datasets to learn general representations and cross-modal interactions. Second, the adaptation stage fine-tunes the pre-trained MLLM on task-specific datasets, enabling it to specialize its knowledge and skills for particular applications. Finally, the evaluation stage assesses the adapted MLLM’s performance by processing multimodal inputs and generating appropriate outputs, ensuring its effectiveness in solving real-world problems (Li et al., 2022b; Liu et al., 2024b, 2023c; Zhang et al., 2024c).
From a data-centric perspective, the performance of MLLMs heavily relies on the quality and diversity of the multimodal datasets used throughout their development stages (Liu et al., 2024b, 2023c). In the pre-training stage, large-scale datasets encompassing multiple modalities are essential for learning robust and generalizable representations (McKinzie et al., 2024). During the adaptation stage, task-specific datasets with high-quality annotations are crucial for effective fine-tuning (Liu et al., 2023c). Lastly, in the evaluation stage, MLLMs must be assessed on their ability to handle diverse and potentially noisy multimodal inputs encountered in real-world scenarios (Zhang et al., 2024c).
@@
MLLMs represent a significant advancement in AI by extending the capabilities of traditional LLMs to process and analyze multiple modalities. By leveraging large-scale multimodal datasets and specialized training techniques, MLLMs can solve complex real-world problems that require understanding and reasoning across different types of information. As research in this field progresses, we can expect MLLMs to play an increasingly important role in various domains, from robotics and healthcare to education and entertainment.
@@
Data-Centric MLLMsData Colloctingand ProcessingData CollectingSourcesCommonWebpages CommonCrawl Project (Raffel et al., 2020), Self-Crawled Data: General Approches (Sharma et al., 2018; Changpinyo et al., 2021), Focused Approaches (Wikipedia (Gao et al., 2020; Srinivasan et al., 2021; Schamoni et al., 2018), Flickr (Young et al., 2014), FreeSound (Mei et al., 2023; Fonseca et al., 2022)), Querying search engines-based approaches (Wu et al., 2017; Gu et al., 2022) Social Media Stack Exchange (Soboleva et al., 2023; Chen et al., 2023g), Reddit (Baumgartner et al., 2020; Desai et al., 2021; Overbay et al., 2023), YouTube (Miech et al., 2019; Zhou et al., 2017), X (Barbieri et al., 2020) Academic Papers ArXiv (Computer, 2023; Gao et al., 2020; Liu et al., 2023d), S2ORC (Lo et al., 2020; Soldaini and Lo, 2023; Soldaini et al., 2024) Books Project Gutenberg (Rae et al., 2019), Smashwords (Zhu et al., 2015), Bibliotik (Gao et al., 2020), Amazon.com (Mishra et al., 2019), Old Photo Books (Okamoto et al., 2023) Codes BigQuery (Nijkamp et al., 2022), GitHub (Kocetkov et al., 2022; Nijkamp et al., 2022; Gao et al., 2020), The Stack (Li et al., 2023a) ProfessionalSources Legal Domain: FreeLaw (Gao et al., 2020); Math Domain: Khan Academy, DeepMind (Saxton et al., 2018); Medical Domain: Online Medical Websites (Li et al., 2023j; Ben Abacha and Demner-Fushman, 2019), Knowledge Bases (Li et al., 2023j; Welbl et al., 2018), In-Hospital Database Systems (Johnson et al., 2023, 2019; Bustos et al., 2020); Financial Domain: Financial Platforms and Forums (Ok, 2023; Lu et al., 2023b) Data ProcessingMethodsFiltering Textual Filtering (Joulin et al., 2016; Grave et al., 2018; Raffel et al., 2020, 2020; Yuan et al., 2021; Gunasekar et al., 2023), Image-Level Filtering (Zhu et al., 2024; Guo et al., 2021; Yang et al., 2022), Video-Level Filtering (Wang et al., 2023a; Blattmann et al., 2023; Farnebäck, 2003; Baek et al., 2019) Deduplication Exact Duplication (Raffel et al., 2020; Suárez et al., 2019; Lee et al., 2022), Approximate Duplication (Computer, 2023; Lee et al., 2022; Penedo et al., 2023), Sentence-Level and Document-Level Semantic Duplication (Reimers and Gurevych, 2019; Song et al., 2020; Silcock et al., 2022; Abbas et al., 2023) DataEnhancement Enhance X Modality Data (Li et al., 2023k; Luo et al., 2024; Wei et al., 2023), Enhance Text Data (Nguyen et al., 2023; Fan et al., 2024; Lai et al., 2023a; Chen et al., 2023e) Commonly-Used DatasetsImage datasets Image-Caption Datasets (Schuhmann et al., 2021, 2022, 2022), Content Descriptive Image-Caption Dataset (Chen et al., 2015; Young et al., 2014; Chen et al., 2022), Interleaved Image-Text Documents (Alayrac et al., 2022; Zhu et al., 2024; Laurençon et al., 2024), Visual Question Answer Image-Caption Datasets (Goyal et al., 2019; Zhu et al., 2016; Biten et al., 2019) Video datasets Video Captioning Datasets (Xu et al., 2016; Bain et al., 2021; Alayrac et al., 2022), Video Question Answering Datasets (Chen and Dolan, 2011; Xu et al., 2016; Jang et al., 2017), Interleaved Video-Text Dataset (Sun et al., 2023b) Audio datasets Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022c), Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) 3D datasets 3D Captioning Datasets (Dai et al., 2017; Armeni et al., 2017; Zheng et al., 2020a) Data-CentricPre-trainingDomain Mixture Pure Text Domain Mixture (Xie et al., 2024a, b; Du et al., 2022), Multimodal Domain Mixture (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a) Modality Mixture Image Caption Data Mixture (McKinzie et al., 2024; Lai et al., 2023b), Image and Video Caption Data Mixture (Lin et al., 2023b; Zhang et al., 2023; Luo et al., 2023; Jin et al., 2023; Li et al., 2023g; Han et al., 2023; Li et al., 2023c, h; Wang et al., 2023a; Chen et al., 2023c) Quality Selection Active Learning-Based Selection (Xu et al., 2023a), Selection Before Training (Gadre et al., 2024), Distribution-Agnostic Selection (Gadre et al., 2024; Wang et al., 2024d; Mahmoud et al., 2024), Distribution-Aware Selection (Wang et al., 2024a; Gadre et al., 2024) Data-Centric AdaptationSupervisedFinetuningData Processing SFT Data Processing (Zhu et al., 2023; Chen et al., 2023e) Data Generation Captioning Instruction-Response Datasets Generation (Zhu et al., 2023; Xu et al., 2024; OpenAI, 2023), Question Answer Instruction-Response Datasets Generation (Liu et al., 2023c; Singh et al., 2019; Li et al., 2023l), Reasoning Instruction-Response Datasets Generation (Chen et al., 2023j, h, m), Other Instruction-Response Datasets Generation (Li et al., 2023l; Panagopoulou et al., 2023; Dai et al., 2024) Data Selection Coreset-Based Selection (Chen et al., 2023k; Das and Khetan, 2023; Sorscher et al., 2022), LLMs-Based Selection (Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e), gradient-Based Selection (Paul et al., 2021; Attendu and Corbeil, 2023; Xia et al., 2024), Self-Instruction-Based Selection Methods (Li et al., 2023m, d; Kung et al., 2023; Liu et al., 2024c) Alignment Human Preference Alignment: Pure Text (Ouyang et al., 2022; Li et al., 2023g), Multimodal (Liu et al., 2023b; Sun et al., 2023a; Chen et al., 2023i; Ouyang et al., 2022) Data-CentricEvaluationData Evaluation Dataset Diversity (Heusel et al., 2017; Friedman and Dieng, 2023); Dataset Quality (Honovich et al., 2022a; Rohrbach et al., 2018; Jing et al., 2023); Dataset Similarity (Sun and Saenko, 2016; Jiang et al., 2022; Pillutla et al., 2023) Evaluation DatasetsCaptioningTasks Image Captioning Datasets (Goyal et al., 2019; Hudson and Manning, 2019; Marino et al., 2019), Video Captioning Datasets (Xu et al., 2017; Chen and Dolan, 2011; Yu et al., 2019), Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022c) QuestionAnswering Tasks Image Question Answering Datasets (Chen et al., 2015; Young et al., 2014; Karpathy and Fei-Fei, 2015), Video Question Answering Datasets (Chen and Dolan, 2011; Venugopalan et al., 2015; Xu et al., 2016), Audio Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) Perception &Reasoning Tasks Perception and Reasoning Datasets (Singh et al., 2019; Masry et al., 2022; Kembhavi et al., 2016), Spatial Reasoning Datasets (Liu et al., 2023b; Kazemzadeh et al., 2014; Mao et al., 2016) Other NoteworthyTasks Classification Datasets (Kiela et al., 2020; Deng et al., 2009; Piczak, 2015; Wu et al., 2015; Goyal et al., 2017), Multimodal Dialogue Datasets (Bai et al., 2023b)
@@
2.3. Data-Centric AI and Why Data-Centric LLMs

### v2

@@ Abstract.
Multimodal large language models (MLLMs) enhance the capabilities of standard large language models by integrating and processing data from multiple modalities, including text, vision, audio, video, and 3D environments. Data plays a pivotal role in the development and refinement of these models. In this survey, we comprehensively review the literature on MLLMs from a data-centric perspective. Specifically, we explore methods for preparing multimodal data during the pretraining and adaptation phases of MLLMs. Additionally, we analyze the evaluation methods for the datasets and review the benchmarks for evaluating MLLMs. Our survey also outlines potential future research directions. This work aims to provide researchers with a detailed understanding of the data-driven aspects of MLLMs, fostering further exploration and innovation in this field.
Large Language Models; Generative Models; Multimodal Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Computer vision
@@
In recent years, we have witnessed rapid advancement of large language models (LLMs) and multimodal large language models (MLLMs) (Zhao et al., 2023c; Wu et al., 2023b). MLLMs, such as GPT-4 (OpenAI, 2023), Flamingo (Alayrac et al., 2022), LLaVA (Liu et al., 2024b), BLIP2 (Li et al., 2023e), Cambrian-1 (Tong et al., 2024) and X-InstructBLIP (Panagopoulou et al., 2023), integrate multiple modality information, demonstrating impressive comprehension and generation capabilities. These models achieve competitive performance in traditional multimodal tasks, such as visual recognition (Zhang et al., 2024a), video understanding (Xu et al., 2021; Tang et al., 2023), speech recognition (Min and Wang, 2023) and 3D understanding (Guo et al., 2023; Hong et al., 2023). Moreover, their excellent language understanding capacity enables strong performance in text-rich tasks, such as question answering (Hu et al., 2024), multi-dialog conversation and logical reasoning (Xu et al., 2024; Li et al., 2023k).
Most existing MLLMs focus on modifying model architecture to explore the use of information from modalities (Xu et al., 2024; Panagopoulou et al., 2023; Alayrac et al., 2022). While model effectiveness is crucial, data also significantly impacts the success of MLLMs. For example, Hoffmann et al. (2022) shows that in order to scale up models, it is necessary to increase the scale of the training data. Beyond data volume, data quality is equally important. Previous research (Sorscher et al., 2022) indicates that carefully curated datasets can enable smaller models to achieve comparable performance to larger ones. However, comprehensive studies on data curation and utilization for MLLMs are still lacking. Therefore, this study aims to provide a comprehensive understanding of MLLMs from a data-centric perspective.
In contrast to model-centric approaches that prioritize architectural enhancements while relying on fixed datasets, data-centric perspectives emphasize the comprehensive impact of the training corpus datasets on the model performance. Within the scope of data-centric MLLMs, our focus lies on leveraging the heterogeneous nature of data modalities, enhancing data structure, increasing data quantity, and elevating data quality to augment MLLMs (Zha et al., 2023). Our survey discusses three key questions from a data-centric perspective at different stages of MLLMs:
  • Q1: How to collect, select, and manage data for MLLMs? The substantial data volume requirements and the heterogeneity of multimodal data pose challenges in gathering, selecting, and effectively managing data for model training. Different training stages of MLLMs also lead to varying data type requirements.
@@
Unique contribution of this survey. Several existing works have focused on LLMs (Zhao et al., 2023c; Hadi et al., 2023; Naveed et al., 2023) and MLLMs (Wu et al., 2023b; Zhang et al., 2024c) from a model-centric perspective, but lack an in-depth analysis from the data-centric perspective. Recently, some work has started to focus on data preparation for LLMs, such as data management methods (Wang et al., 2023b), data selection methods (Albalak et al., 2024), and comprehensive reviews of LLM datasets (Liu et al., 2024a). However, such work primarily focuses on data management and selection methods for text-only LLMs and does not provide a detailed analysis of the data processing pipeline for MLLMs. The most closely related work to ours is data-centric artificial intelligence (DCAI) (Zha et al., 2023; Jarrahi et al., 2022; Jakubik et al., 2022; Polyzotis and Zaharia, 2021; Whang et al., 2023), which also focuses on data-centric views of AI research but does not specifically analyze LLMs and MLLMs.
@@
In this survey, we reviewed the literature on the advances of MLLMs from data-centric perspectives. We provide researchers and developers with a general and comprehensive understanding of the latest developments in the field of MLLM data side. The key contributions of this survey are summarized as follows:
@@
  • Data evaluation methods and evaluation benchmarks. We outline commonly used methods to evaluate datasets as well as MLLMs evaluation benchmarks from a data-centric perspective.
@@
The rest of this survey is organized as follows: Section 2 introduces the preliminaries for LLMs and MLLMs, and discusses the motivations to analyze them from a data-centric perspective. Sections 3 to 5 summarize the main stages of collecting, processing, and selecting data for MLLMs training. Section 6 enumerates the data evaluation methods and existing evaluation datasets for MLLMs according to evaluation tasks. Section 7 discusses open issues and highlights several future research directions in this field. Finally, we conclude the survey in Section 8.
@@ 2.1. Large Language Models
A significant breakthrough came with the introduction of pre-trained language models, which aimed to capture general language representations by training on vast amounts of unlabeled text data (Kenton and Toutanova, 2019; Radford et al., [n. d.]). Building upon these advancements, LLMs have taken the concept of pre-training to an unprecedented scale, with models containing billions of parameters and trained on massive corpora spanning hundreds of gigabytes to terabytes of text data (Brown et al., 2020; Raffel et al., 2020).
@@
LLMs exhibit several notable properties that distinguish them from previous language models. One fundamental characteristic is the scaling law, which describes how the performance of large language models improves as they scale in terms of model size, training data, and computational resources (Kaplan et al., 2020; Hoffmann et al., 2022). This power-law relationship suggests that larger models trained on more data tend to capture more complex patterns and generalize better to new tasks. Another intriguing property of LLMs is the emergence of abilities that were not explicitly trained for, often referred to as emergent abilities (Wei et al., 2022). This ability suggests that LLMs can capture and leverage complex linguistic patterns and knowledge from the pre-training data, enabling them to perform tasks beyond their original training objective. Furthermore, LLMs have demonstrated the capability of in-context learning (Brown et al., 2020), where they can perform tasks based on a few examples provided in the input prompt, without the need for explicit fine-tuning. This highlights the models’ ability to rapidly adapt to new tasks and generalize from limited examples.
@@
The evolution of LLMs typically involves three key stages: training, adaptation, and evaluation (Zhao et al., 2023c). The training stage focuses on learning general language representations from large-scale unlabeled corpora, capturing the underlying patterns and structures of natural language. The adaptation stage involves guiding the pre-trained model to specific tasks and human preference, often through supervised finetuning and human preference alignment (Raffel et al., 2020; Ouyang et al., 2022). This step is crucial for optimizing the model’s performance on the target application and ensuring its outputs are aligned with the desired objectives and human preference. Finally, the evaluation stage involves assessing the performance of the adapted model through various metrics and benchmarks to ensure that it meets the required standards and performs effectively in real-world scenarios.
@@ 2.2. Multimodal Large Language Models
Multimodal large language models (MLLMs) extend the capabilities of traditional LLMs by leveraging their comprehension, generation, and task-solving abilities across different modalities, including natural language (text and audio) and visual information (videos, images, and 3D models). The development of MLLMs is driven by the need to address real-world problems that often involve multiple modalities. For instance, in robotics, MLLM can process visual input from cameras, interpret natural language instructions, and analyze structured data from sensors to perform complex tasks (Zeng et al., 2023). Similarly, in healthcare, MLLMs can analyze medical images, process electronic health records, and interpret patient-doctor conversations to aid in diagnosis and treatment planning.
Multimodal large language models (MLLMs) typically consist of three primary components: a modality encoder, a projector, and a large language model (LLM). The modality encoder is responsible for encoding data from different modalities, the projector aligns the data from various modalities with the LLM, and the LLM serves as the backbone, providing comprehension capabilities. By integrating LLMs with multimodal encoders and projectors, MLLMs can process various types of information, demonstrating strong understanding and analytical capabilities to address downstream tasks across various modalities (Wu et al., 2023b).
@@
MLLMs represent a significant advancement in AI by extending the capabilities of traditional LLMs to process and analyze multiple modalities. MLLMs can solve complex real-world problems that require understanding and reasoning across different types of information. As research in this field progresses, we can expect MLLMs to play an increasingly important role in various domains, from robotics and healthcare to education and entertainment.
@@
Data-Centric MLLMsData Collectingand ProcessingData CollectingSourcesCommonWebpages CommonCrawl Project (Raffel et al., 2020), Self-Crawled Data: General Approches (Sharma et al., 2018; Changpinyo et al., 2021), Focused Approaches (Wikipedia (Gao et al., 2020; Srinivasan et al., 2021; Schamoni et al., 2018), Flickr (Young et al., 2014), FreeSound (Mei et al., 2023; Fonseca et al., 2022)), Querying search engines-based approaches (Wu et al., 2017; Gu et al., 2022) Social Media Stack Exchange (Soboleva et al., 2023; Chen et al., 2023g), Reddit (Baumgartner et al., 2020; Desai et al., 2021; Overbay et al., 2023), YouTube (Miech et al., 2019; Zhou et al., 2017), X (Barbieri et al., 2020) Academic Papers ArXiv (Computer, 2023; Gao et al., 2020; Liu et al., 2023d), S2ORC (Lo et al., 2020; Soldaini and Lo, 2023; Soldaini et al., 2024) Books Project Gutenberg (Rae et al., 2019), Smashwords (Zhu et al., 2015), Bibliotik (Gao et al., 2020), Amazon.com (Mishra et al., 2019), Old Photo Books (Okamoto et al., 2023) Codes BigQuery (Nijkamp et al., 2022), GitHub (Kocetkov et al., 2022; Nijkamp et al., 2022; Gao et al., 2020), The Stack (Li et al., 2023a) ProfessionalSources Legal Domain: FreeLaw (Gao et al., 2020); Math Domain: Khan Academy, DeepMind (Saxton et al., 2018); Medical Domain: Online Medical Websites (Li et al., 2023i; Ben Abacha and Demner-Fushman, 2019), Knowledge Bases (Li et al., 2023i; Welbl et al., 2018), In-Hospital Database Systems (Johnson et al., 2023, 2019; Bustos et al., 2020); Financial Domain: Financial Platforms and Forums (Ok, 2023; Lu et al., 2023b) Data ProcessingMethodsFiltering Textual Filtering (Joulin et al., 2016; Grave et al., 2018; Raffel et al., 2020, 2020; Yuan et al., 2021; Gunasekar et al., 2023), Image-Level Filtering (Zhu et al., 2024; Guo et al., 2021; Yang et al., 2022), Video-Level Filtering (Wang et al., 2023a; Blattmann et al., 2023; Farnebäck, 2003; Baek et al., 2019) Deduplication Exact Duplication (Raffel et al., 2020; Suárez et al., 2019; Lee et al., 2022), Approximate Duplication (Computer, 2023; Lee et al., 2022; Penedo et al., 2023), Sentence-Level and Document-Level Semantic Duplication (Reimers and Gurevych, 2019; Song et al., 2020; Silcock et al., 2022; Abbas et al., 2023) DataEnhancement Enhance X Modality Data (Li et al., 2023j; Luo et al., 2024; Wei et al., 2023), Enhance Text Data (Nguyen et al., 2023; Fan et al., 2024; Lai et al., 2023a; Chen et al., 2023e) Commonly-Used DatasetsImage datasets Image-Caption Datasets (Schuhmann et al., 2021, 2022, 2022), Content Descriptive Image-Caption Dataset (Chen et al., 2015; Young et al., 2014; Chen et al., 2022), Interleaved Image-Text Documents (Alayrac et al., 2022; Zhu et al., 2024; Laurençon et al., 2024), Visual Question Answer Image-Caption Datasets (Goyal et al., 2019; Zhu et al., 2016; Biten et al., 2019) Video datasets Video Captioning Datasets (Xu et al., 2016; Bain et al., 2021; Alayrac et al., 2022), Video Question Answering Datasets (Chen and Dolan, 2011; Xu et al., 2016; Jang et al., 2017), Interleaved Video-Text Dataset (Sun et al., 2023b) Audio datasets Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022b), Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) 3D datasets 3D Captioning Datasets (Dai et al., 2017; Armeni et al., 2017; Zheng et al., 2020) Data-CentricPre-trainingDomain Mixture Pure Text Domain Mixture (Xie et al., 2024a, b; Du et al., 2022), Multimodal Domain Mixture (Xu et al., 2023c; Chen et al., 2024b; Song et al., 2023; Chen et al., 2023l; Wang et al., 2023a) Modality Mixture Image Caption Data Mixture (McKinzie et al., 2024; Lai et al., 2023b), Image and Video Caption Data Mixture (Lin et al., 2023b; Zhang et al., 2023; Luo et al., 2023; Jin et al., 2023; Li et al., 2023f; Han et al., 2023; Li et al., 2023c, g; Wang et al., 2023a; Chen et al., 2023c) Quality Selection Active Learning-Based Selection (Xu et al., 2023a), Selection Before Training (Gadre et al., 2024), Distribution-Agnostic Selection (Gadre et al., 2024; Wang et al., 2024c; Mahmoud et al., 2024), Distribution-Aware Selection (Wang et al., 2024a; Gadre et al., 2024) Data-Centric AdaptationSupervisedFinetuningData Processing SFT Data Processing (Zhu et al., 2023; Chen et al., 2023e) Data Generation Captioning Instruction-Response Datasets Generation (Zhu et al., 2023; Xu et al., 2024; OpenAI, 2023), Question Answer Instruction-Response Datasets Generation (Liu et al., 2023c; Singh et al., 2019; Li et al., 2023k), Reasoning Instruction-Response Datasets Generation (Chen et al., 2023j, h, m), Other Instruction-Response Datasets Generation (Li et al., 2023k; Panagopoulou et al., 2023; Dai et al., 2024) Data Selection Coreset-Based Selection (Chen et al., 2023k; Das and Khetan, 2023; Sorscher et al., 2022), LLMs-Based Selection (Chen et al., 2023f; Xu et al., 2023b; Liu et al., 2023e), gradient-Based Selection (Paul et al., 2021; Attendu and Corbeil, 2023; Xia et al., 2024), Self-Instruction-Based Selection Methods (Li et al., 2023l, d; Kung et al., 2023; Liu et al., 2024c) Alignment Human Preference Alignment: Pure Text (Ouyang et al., 2022; Li et al., 2023f), Multimodal (Liu et al., 2023b; Sun et al., 2023a; Chen et al., 2023i; Ouyang et al., 2022) Data-CentricEvaluationData Evaluation Dataset Diversity (Heusel et al., 2017; Friedman and Dieng, 2023); Dataset Quality (Honovich et al., 2022a; Rohrbach et al., 2018; Jing et al., 2023); Dataset Similarity (Sun and Saenko, 2016; Jiang et al., 2022; Pillutla et al., 2023) Evaluation DatasetsCaptioningTasks Image Captioning Datasets (Goyal et al., 2019; Hudson and Manning, 2019; Marino et al., 2019), Video Captioning Datasets (Xu et al., 2017; Chen and Dolan, 2011; Yu et al., 2019), Audio Captioning Datasets (Lipping et al., 2022; Li et al., 2022b) QuestionAnswering Tasks Image Question Answering Datasets (Chen et al., 2015; Young et al., 2014; Karpathy and Fei-Fei, 2015), Video Question Answering Datasets (Chen and Dolan, 2011; Venugopalan et al., 2015; Xu et al., 2016), Audio Question Answering Datasets (Kim et al., 2019; Drossos et al., 2020) Perception &Reasoning Tasks Perception and Reasoning Datasets (Singh et al., 2019; Masry et al., 2022; Kembhavi et al., 2016), Spatial Reasoning Datasets (Liu et al., 2023b; Kazemzadeh et al., 2014; Mao et al., 2016) Other NoteworthyTasks Classification Datasets (Kiela et al., 2020; Deng et al., 2009; Piczak, 2015; Wu et al., 2015; Goyal et al., 2017), Multimodal Dialogue Datasets (Bai et al., 2023b)
@@
2.3. Data-Centric AI and Why Data-Centric MLLMs